OpenStudy (anonymous):

Use the function to answer the question. f(x) = 10x - 5 What is the value of f^–1 (–4)? (Points : 4) -45 -35 0.01 0.1

OpenStudy (anonymous):

Best way is just take the function's name called 10x - 5 and set it to -4. Reason is beacuse you want the value of the inverse function.

OpenStudy (anonymous):

I agree with easy, any more info would hand you the answer

OpenStudy (anonymous):

So would it be the 2nd one...?

OpenStudy (anonymous):

what is 10(-4) -5 ?

OpenStudy (anonymous):

-35 right...?

OpenStudy (anonymous):

Ehuman...you dont evaluate 10x - 5 at x = 4, I said take 10x - 5 = 4.

OpenStudy (anonymous):

what is 10*-4?

OpenStudy (anonymous):

-40

OpenStudy (anonymous):

i never said that he needs to evaluate the function at x =-4

OpenStudy (anonymous):

what is -40 - 5?

OpenStudy (anonymous):

-45.... OH

OpenStudy (anonymous):

You are all making an error. You want the INVERSE at - 4, you do NOT want the value of the actual function at x = -4.

OpenStudy (anonymous):

AHHH!!!! SO CONFUSED!! XD

OpenStudy (anonymous):

oh my bad, cart before the horse, 2 lashings with a wet noodle, walking away from the pc for a while lol.

OpenStudy (anonymous):

listen to easy's first advise

OpenStudy (anonymous):

10x - 5 = -4 solve

OpenStudy (anonymous):

10x - 5 - -4 10x = 1 x = 1/10 or 0.1